OpenStudy (anonymous):

14.925 in fractional notation

OpenStudy (anonymous):

\[14.925 = \frac{14.925}{1} = \frac{ 149.25}{10} = \frac{1492.5}{100}\]

OpenStudy (anonymous):

597/40

OpenStudy (anonymous):

keep going until the numerator does not have a decimal, then simplify the fraction